# Robust Adaptive Discrete-Time Control Barrier Certificate

### v1

@@ 2.3 Robust safety certificate
If a continuous function $B:\mathcal{X}\to\mathbb{R}$ is a robust DT-CBF for (6) on $\mathcal{S}_{B}$, then $\mathcal{U}_{\mathrm{rs}}(x;\theta^{\ast})$ is non-empty for all $x\in\mathcal{X}$. In addition, if a state-feedback control policy $\pi:\mathbb{R}^{n}\to\mathbb{R}^{m}$ satisfies $\pi(x)\in\mathcal{U}_{\mathrm{rs}}(x;\theta^{\ast})$, $\forall x\in\mathcal{X}$, then the closed-loop system $x_{t+1}=f(x_{t},\pi(x_{t});\theta^{\ast})+w_{t}$ is robustly safe w.r.t. $\mathcal{S}_{B}$ against disturbances $w_{t}$.
Remark 2.1.
Contrary to the existing results on robust DT-CBFs with (unbounded) stochastic disturbances [12], [13], this paper focuses on bounded disturbances that can be stochastic [28] or adversarial [1]. Accordingly, Definitions 1 and 2 and Theorem 1 are tailored for robust safety against bounded disturbances only, assuming the true parameter $\theta^{\ast}$ is known to the controller.
Since knowing $\theta^{\ast}$ is a stringent requirement in general, (10) cannot be used to synthesize safe controllers when $\theta^{\ast}$ is unknown. In such cases, a straightforward remedy is to additionally robustify against maximized model mismatch and use a constant nominal $\hat{\theta}$ to build the CBF condition. Accordingly, similar to Theorem 1, a corresponding worst-case robust safety guarantee can be established. However, controllers designed based on such worst-case safety condition may not respond to dangerous system behaviors as fast as its counterpart that knows the true model, and more conservative inputs are then needed to guarantee safety. Considering that parametric uncertainty is epistemic in nature [43] and can be reduced or even eliminated through the acquisition of data, it is thus beneficial to develop an adaptive barrier certificate that can be used for safe control synthesis.
3 Robust adaptive safety
In this section, we introduce robust adaptive safety guarantees with general online parameter adaptation laws. More specifically, a robust adaptive CBC is formulated that only relies on an instantaneous parameter estimate and the associated adaptation law. This proposed certificate ensures robust safety against both parametric model mismatch and disturbances with reduced conservatism due to the adaptation mechanism.
@@ 3.1 Online parameter estimation
At each time step $t$, consider a parameter estimation module $\mathtt{EST}$ that can provide a polytopic set estimate $\Theta_{t}\subseteq\Theta$ such that $\theta^{\ast}\in\Theta_{t}$ and a point estimate $\hat{\theta}_{t}\in\Theta_{t}$. Following (4) and (5), the instantaneous estimation error and its associated error bound are defined as follows:
@@
$\varepsilon_{\theta,t}(p):=\max_{\theta\in\Theta_{t}}\|\theta-\hat{\theta}_{t}\|_{p}.\vskip-5.69046pt$ (11b)
@@
where $\mathcal{T}_{\mathtt{ADP},t}(\cdot)$ is time-varying since the closed-loop input-state data is collected online and the set estimate $\Theta_{t}$ will also be used to determine the point estimate. Finally, the point-estimate increment is defined by
$\delta_{\theta,t}:=\hat{\theta}_{t+1}-\hat{\theta}_{t}=\mathcal{T}_{\mathtt{ADP},t}(\hat{\theta}_{t})-\hat{\theta}_{t},\vskip-5.69046pt$ (13)
which will be used to build the barrier certificate.
3.2 Robust adaptive control barrier certificate
Given a baseline continuous barrier candidate $B:\mathcal{X}\to\mathbb{R}$ that encodes safety constraints and that defines a safe set $\mathcal{S}_{B}$ as (7), we define the corresponding robust adaptive barrier function by
$B_{\mathrm{ra},t}(x):=B(x)-\frac{1}{2}e_{\theta,t}^{\top}\Gamma^{-1}e_{\theta,t},\vskip-5.69046pt$ (14)
@@ 3.2 Robust adaptive control barrier certificate
$\mathcal{S}_{\mathrm{ra},t}:=\{x\in\mathcal{X}\mid B_{\mathrm{ra},t}(x)\geq 0\},\vskip-5.69046pt$ (15)
and its boundary $\partial\mathcal{S}_{\mathrm{ra},t}$ and interior $\mathrm{int}(\mathcal{S}_{\mathrm{ra},t})$ are defined similar to (8a) and (8b), respectively. Besides, it is noted that (14) implies that $\mathcal{S}_{\mathrm{ra},t}\subseteq\mathcal{S}$, and $\mathcal{S}_{\mathrm{ra},t}=\mathcal{S}$ if and only if $\hat{\theta}_{t}=\theta^{\ast}$. Since $\mathcal{S}_{\mathrm{ra},t}$ is time-varying, a timed version of the robust safety in Definition 1 is given as follows:
@@
Given $T\in\overline{\mathbb{N}}$, a sequence of sets $\{\mathcal{X}_{\mathrm{s},t}\}^{T}_{t=0}$ satisfying $x_{0}\in\mathcal{X}_{\mathrm{s},0}$ and $\forall t\in\mathbb{I}_{[0:T]}$, $\mathcal{X}_{\mathrm{s},t}\subseteq\mathcal{X}$ is sequentially positively invariant for (6) if $\forall t\in\mathbb{I}_{[0:T-1]}$, $x_{t}\in\mathcal{X}_{\mathrm{s},t}$ implies that $x_{t+1}\in\mathcal{X}_{\mathrm{s},t+1}$ for all $w_{t}\in\mathcal{W}$. In this case, the system (6) is said to be robustly safe w.r.t. $\{\mathcal{X}_{\mathrm{s},t}\}^{T}_{t=0}$. In addition, if $\mathcal{X}_{\mathrm{s},t}=\mathcal{X}_{\mathrm{s}}^{\ast}$, $\forall t\in\mathbb{I}_{[0:T]}$, the system (6) is said to be robustly safe w.r.t. $\mathcal{X}_{\mathrm{s}}^{\ast}$ over $\mathbb{I}_{[0:T-1]}$.
@@
In this work, we consider a general barrier function $B(\cdot)$, instead of a parameter-dependent barrier function $B(\cdot;\theta)$ as in [25], [38]. In fact, model-free CBFs that are derived solely based on the safety requirements have been used in many real-world safe control applications, e.g., ACC [25], [38], aircraft pitch control [25], and aerial robotics [48]. This framework also avoids repeatedly constructing barrier functions $B(\cdot;\hat{\theta}_{t})$ online based on $\hat{\theta}_{t}$, which can be computationally expensive [35].
Since $\theta^{\ast}$ is unknown, $e_{\theta,t}$ cannot be computed, and thus $\mathcal{S}_{\mathrm{ra},t}$ cannot be explicitly characterized. However, $\mathcal{S}_{\mathrm{ra},t}$ only serves as a fictitious set that will be used to realize safety characterized by $B(\cdot)$, as will be manifested in Theorem 2. Next, we present the robust adaptive CBC, which certifies $B(\cdot)$ as a barrier function for a system in the presence of bounded disturbances and parameteric uncertainty with adaptation. In addition, this CBC can be used in the safety filter framework [4].
Definition 3.3 (Robust adaptive CBC).
Let Assumption 1 hold. Given $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, a continuous function $B:\mathcal{X}\to\mathbb{R}$ is a robust adaptive DT-CBF for (6) and $\mathtt{EST}$ on $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ if there exists $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$ such that, for all $x\in\mathcal{X}$, there exists $u\in\mathcal{U}$ satisfying
$B\big{(}f(x,u;\hat{\theta}_{t})\big{)}-B(x)-L_{B}\overline{w}-E_{\theta,t}(x)\\ \geq-\alpha\left(B(x)-\frac{(\varepsilon_{\theta,t}(p))^{2}}{2\lambda_{\min}(\Gamma)}\right),\forall t\in\mathbb{I}_{[0:T-1]},\vskip-5.69046pt$ (16)
where $E_{\theta,t}(x)\hskip-2.27626pt=\hskip-2.27626pt\left(L_{B}\|\bm{\phi}(x)\|+\frac{\|\delta_{\theta,t}\|}{\lambda_{\min}(\Gamma)}\right)\varepsilon_{\theta,t}(p)+\frac{\|\delta_{\theta,t}\|^{2}}{2\lambda_{\min}(\Gamma)}$ with $L_{B}$ being the Lipschitz constant in (9) and $\delta_{\theta,t}$ characterized by the estimation module $\mathtt{EST}$. In this case, the condition (16) is called the robust adaptive CBC against disturbances and parametric model mismatch.
The certificate (16) also defines a set of robust adaptive safe control inputs for $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ as
[…]
where the parametric dependency on $\hat{\theta}_{t}$ indicates that characterizing the robust adaptive safe control set only requires the online estimate $\hat{\theta}_{t}$ without knowing the true $\theta^{\ast}$. In essence, $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$ is time-varying and it depends on the estimation module $\mathtt{EST}$. Robust adaptive safety is summarized in the following theorem:
Theorem 2 (Robust adaptive safety).
Let Assumption 1 hold and assume $x_{0}\in\mathcal{S}_{\mathrm{ra},0}$. Given $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, if $B$ is a robust adaptive DT-CBF for the system (6) and $\mathtt{EST}$ on $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$, then the set $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$ is non-empty for all $x\in\mathcal{X}$ and all $t\in\mathbb{I}_{[0,T-1]}$. In addition, if a sequence of time-varying state-feedback control policies $\pi_{t}:\mathcal{X}\to\mathbb{R}^{m},t\in\mathbb{I}_{[0,T-1]}$ satisfies $\pi_{t}(x)\in\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$ for all $x\in\mathcal{X}$, then the time-varying closed-loop system
$x_{t+1}=f(x_{t},\pi_{t}(x_{t});\theta^{\ast})+w_{t}\vskip-5.69046pt$ (18)
is robustly safe w.r.t. $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ and thus robustly safe w.r.t. $\mathcal{S}_{B}$ over $\mathbb{I}_{[0:T]}$ (cf. Definition 3.1) against both parametric model mismatch and disturbances.
The proof of Theorem 2 is given in Appendix A. It is worth noting that the certificate (16) requires only $\hat{\theta}_{t}$ and $\varepsilon_{\theta,t}(p)$, without needing knowledge of $\theta^{\ast}$. This advantage makes (16) more readily applicable for CBF-based controller synthesis. Specifically, given a nominal state-feedback policy $\pi_{\mathrm{nom},t}:\mathcal{X}\to\mathbb{R}^{m}$ at time step $t$, a robust adaptive safe policy $\pi_{\mathrm{ras},t}:\mathcal{X}\to\mathcal{U}$ is implicitly given by the following safety filter:
$\pi_{\mathrm{ras},t}(x;p)=\operatorname*{arg\,min}_{\nu\in\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)}\|\nu-\pi_{\mathrm{nom},t}(x)\|.\vskip-5.69046pt$ (19)
Under $\pi_{\mathrm{ras},t}(\cdot)$, the closed-loop system becomes
$x_{t+1}=f(x_{t},\pi_{\mathrm{ras},t}(x_{t};p);\theta^{\ast})+w_{t}.\vskip-5.69046pt$ (20)

### v2

@@ 2.3 Robust safe control
If the continuous function $B_{\theta^{\ast}}:\mathcal{X}\to\mathbb{R}$ is a robust DT-CBF for (6) on $\mathcal{S}_{B,\theta^{\ast}}$, then $\mathcal{U}_{\mathrm{rs},\theta^{\ast}}(x;\theta^{\ast})$ is non-empty for all $x\in\mathcal{X}$. In addition, if a state-feedback control policy $\pi:\mathcal{X}\to\mathbb{R}^{m}$ satisfies $\pi(x)\in\mathcal{U}_{\mathrm{rs},\theta^{\ast}}(x;\theta^{\ast})$, $\forall x\in\mathcal{X}$, then the closed-loop system $x_{t+1}=f(x_{t},\pi(x_{t});\theta^{\ast})+w_{t}$ is robustly safe w.r.t. $\mathcal{S}_{B,\theta^{\ast}}$ against disturbances $w_{t}$.
Since knowing $\theta^{\ast}$ is generally a stringent requirement, the certificate (8) cannot be directly used to synthesize safe controllers when $\theta^{\ast}$ is unknown. In such cases, a straightforward remedy is to robustify against the worst-case model mismatch and to construct a new inequality condition using a constant nominal $\hat{\theta}$ with additional robustification terms that scale with the estimation error. Accordingly, one can establish a worst-case robust safety guarantee, similar to Theorem 1. However, such worst-case design may react more slowly to dangerous system behaviors compared to those that have access to the true model, and thereby require more conservative control inputs to maintain safety. Since parametric uncertainty is epistemic in nature [40] and can be reduced or even eliminated through online data acquisition, it is desirable to incorporate adaptive safety mechanisms with online model learning to reduce conservatism.
3 Robust adaptive safe control
This section introduces robust adaptive safe control using DT-CBFs with general online parameter estimators that can provide (i) point estimates, (ii) set estimates, and (iii) a uniform bound on the point-estimate increments. More specifically, a sufficient condition used to certify robust adaptive DT-CBFs is provided, and an online input condition is derived to compute safe control inputs in the safety filter framework [12]. The proposed approach can be combined with any nominal controller, enabling modular safe control design.
@@ 3.1 Online parameter estimation
Consider a given parameter set $\Theta$, which in practice is constructed from physical bounds or measurement data [1], typically as a box or polytope enclosing all feasible parameter values. At each time step $t$, consider a parameter estimator $\mathtt{EST}$ that can provide a polytopic set estimate $\Theta_{t}\subseteq\Theta$ such that $\theta^{\ast}\in\Theta_{t}$ and a point estimate $\hat{\theta}_{t}\in\Theta_{t}$. The set $\Theta_{t}$ can be obtained, e.g., through set-membership identification, where the feasible parameter set is updated recursively by intersecting it with the sets consistent with new measurements at each time step. The initial set $\Theta_{0}$ is usually taken as $\Theta$. Following (4) and (5), the instantaneous estimation error and its associated error bound are defined as follows:
@@
$\varepsilon_{\theta,t}(p):=\max_{\theta\in\Theta_{t}}\|\theta-\hat{\theta}_{t}\|_{p}.\vskip-8.5359pt$ (9b)
@@
where $\mathcal{T}_{\mathtt{ADP},t}$ is time-varying since the closed-loop input-state data is collected online and the set estimate $\Theta_{t}$ will also be used to determine the point estimate. Finally, the point-estimate increment is defined by
$\delta_{\theta,t}:=\hat{\theta}_{t+1}-\hat{\theta}_{t}=\mathcal{T}_{\mathtt{ADP},t}(\hat{\theta}_{t})-\hat{\theta}_{t}.\vskip-5.69046pt$ (11)
Moreover, the considered estimator $\mathtt{EST}$ guarantees an uniform upper bound on the increments, i.e.,
$\|\delta_{\theta,t}\|\leq\bar{\delta}_{\theta}.\vskip-5.69046pt$ (12)
Note that such a bound always exists since $\Theta$ is polytopic and therefore compact, implying that $\|\delta_{\theta,t}\|$ always admits a worst-case bound given by $\max_{\theta^{\prime},\theta^{\prime\prime}\in\Theta}\|\theta^{\prime}-\theta^{\prime\prime}\|$.
3.2 Robust adaptive safe control
$\mathcal{X}$$\mathcal{S}_{\text{ra},t}$$\mathcal{S}_{B,\hat{\theta}_{t}}$
(a) Robustification
$\pi_{\text{nom},t}(x)$$\pi_{\text{ras},t}(x;p)$$\mathcal{U}_{\text{ras},t}(x;p)$$\mathcal{U}$
(b) Safety filter
Figure 1: Sketch of (a) adaptively robustified safe sets, and (b) the safety filter projection when the nominal input is unsafe.
Given a CBF $B:\mathcal{X}\times\Theta\to\mathbb{R}$ that encodes safety constraints and that defines a safe set as in (7), we define its corresponding adaptively robustified counterpart by
$B_{\mathrm{ra},t}(x):={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B(x,\hat{\theta}_{t})}-\frac{1}{2}e_{\theta,t}^{\top}\Gamma^{-1}e_{\theta,t},\vskip-5.69046pt$ (13)
@@ 3.2 Robust adaptive safe control
$\mathcal{S}_{\mathrm{ra},t}:=\{x\in\mathcal{X}\mid B_{\mathrm{ra},t}(x)\geq 0\}.\vskip-5.69046pt$ (14)
It is noted that $\mathcal{S}_{\mathrm{ra},t}\subseteq\mathcal{S}_{B,\hat{\theta}_{t}}$ (see Fig. 1(a)), and $\mathcal{S}_{\mathrm{ra},t}=\mathcal{S}_{B,\theta^{\ast}}$ if and only if $\hat{\theta}_{t}=\theta^{\ast}$. Since $\mathcal{S}_{\mathrm{ra},t}$ is time-varying, a timed version of Definition 1 is given as follows:
@@
Given $T\in\overline{\mathbb{N}}$, $\{\mathcal{X}_{\mathrm{s},t}\}^{T}_{t=0}$ satisfying $x_{0}\in\mathcal{X}_{\mathrm{s},0}$ and $\forall t\in\mathbb{I}_{[0:T]}$, $\mathcal{X}_{\mathrm{s},t}\subseteq\mathcal{X}$ is sequentially positively invariant for (6) if $\forall t\in\mathbb{I}_{[0:T-1]}$, $x_{t}\in\mathcal{X}_{\mathrm{s},t}$ implies that $x_{t+1}\in\mathcal{X}_{\mathrm{s},t+1}$ for all $w_{t}\in\mathcal{W}$. In this case, the system (6) is said to be robustly safe w.r.t. $\{\mathcal{X}_{\mathrm{s},t}\}^{T}_{t=0}$. In addition, if $\mathcal{X}_{\mathrm{s},t}\subseteq\mathcal{X}_{\mathrm{s}}^{\ast}$, $\forall t\in\mathbb{I}_{[0:T]}$, the system (6) is said to be robustly safe w.r.t. $\mathcal{X}_{\mathrm{s}}^{\ast}$ over $\mathbb{I}_{[0:T]}$.
@@
Since $\theta^{\ast}$ is unknown, $e_{\theta,t}$ cannot be computed, and thus $\mathcal{S}_{\mathrm{ra},t}$ cannot be explicitly characterized. However, $\mathcal{S}_{\mathrm{ra},t}$ only serves as a fictitious set to realize safety characterized by $B$, as will be manifested in Theorem 2. Next, we present the robust adaptive CBC, which can be used to certify robust adaptive DT-CBFs. Moreover, given such a valid DT-CBF, safe control inputs can be computed in the safety filter framework [3] for systems subject to both bounded disturbances and parametric uncertainty.
Definition 3.2 (Robust adaptive CBF).
Let Assumption 1 hold. Given an estimator $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, a continuous function $B:\mathcal{X}\times\Theta\to\mathbb{R}$ is a robust adaptive DT-CBF for (6) and $\mathtt{EST}$ if there exists $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$ such that, for all $x\in\mathcal{X}$, there exists $u\in\mathcal{U}$ satisfying
$\Delta B{{}^{\star}}(x,u)-L_{B,x}\overline{w}-E^{\star}_{\theta}(x)\\ \geq-\alpha\left(B^{\star}(x)-\frac{(\bar{\varepsilon}_{\theta}(p))^{2}}{2\lambda_{\min}(\Gamma)}\right)\vskip-5.69046pt$ (15)
where $\Gamma$ is given as in (13), $B^{\star}(x):=\min_{\theta\in\Theta}B(x,\theta)$, $\Delta B{{}^{\star}}(x,u)\hskip-2.27626pt:=\hskip-2.27626pt\min_{\theta\in\Theta}\{B(f(x,u;\theta),\theta)\hskip-2.27626pt-\hskip-2.27626ptB(x,\theta)\}$, $\bar{\varepsilon}_{\theta}(p)$ is the worst-case error $p$-norm ($p\in\{1,2\}$) defined as
$\bar{\varepsilon}_{\theta}(p):=\max_{\theta,\theta^{\prime}\in\Theta}\|\theta-\theta^{\prime}\|_{p},\vskip-5.69046pt$ (16)
and $E^{\star}_{\theta}(x)=\left(L_{B,x}\|\bm{\phi}(x)\|+\frac{\bar{\delta}_{\theta}}{\lambda_{\min}(\Gamma)}\right)\bar{\varepsilon}_{\theta}(p)+\frac{\bar{\delta}_{\theta}^{2}}{2\lambda_{\min}(\Gamma)}+L_{B,\theta}\bar{\delta}_{\theta}$, with $L_{B,x}$ and $L_{B,\theta}$ given in Assumption 1 and $\bar{\delta}_{\theta}$ characterized by $\mathtt{EST}$ as in (12). In this case, the condition (15) is called the robust adaptive CBC against disturbances and parametric model mismatch.
Theorem 2 (Robust adaptive safety).
Let Assumption 1 hold and assume $x_{0}\in\mathcal{S}_{\mathrm{ra},0}$. Given an estimator $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, if $B$ is a robust adaptive DT-CBF for (6) and $\mathtt{EST}$ with its associated $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$, then for the robust adaptive online input condition
$B\big(f(x,u;\hat{\theta}_{t}),\hat{\theta}_{t}\big)-B(x,\hat{\theta}_{t})-L_{B,x}\overline{w}\\ -E_{\theta,t}(x)\geq-\alpha\left(B(x,\hat{\theta}_{t})-\frac{(\varepsilon_{\theta,t}(p))^{2}}{2\lambda_{\min}(\Gamma)}\right),\vskip-11.38092pt$ (17)
the set $\mathcal{U}_{\mathrm{ras},t}(x;p)$ defined as
[…]
is non-empty, $\forall x\in\mathcal{X}$ and $t\in\mathbb{I}_{[0,T-1]}$. The term $E_{\theta,t}(x)$ in (17) equals $\left(L_{B,x}\|\bm{\phi}(x)\|+\frac{\|\delta_{\theta,t}\|}{\lambda_{\min}(\Gamma)}\right)\varepsilon_{\theta,t}(p)$ $\frac{\|\delta_{\theta,t}\|^{2}}{2\lambda_{\min}(\Gamma)} +$$L_{B,\theta}\|\delta_{\theta,t}\|$ with $\delta_{\theta,t}$ defined in (11). In addition, if time-varying state-feedback control policies $\pi_{t}:\mathcal{X}\to\mathbb{R}^{m},t\in\mathbb{I}_{[0,T-1]}$ satisfies $\pi_{t}(x)\in\mathcal{U}_{\mathrm{ras},t}(x;p)$ for all $x\in\mathcal{X}$, then the time-varying closed-loop system
$x_{t+1}=f(x_{t},\pi_{t}(x_{t});\theta^{\ast})+w_{t}\vskip-5.69046pt$ (19)
is robustly safe w.r.t. $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ and thus robustly safe w.r.t. $\{\mathcal{S}_{B,\hat{\theta}_{t}}\}^{T}_{t=0}$ (cf. Definition 3.1) against both parametric model mismatch and disturbances.
Verifying a candidate barrier function using the certificate (15) does not require the knowledge of $\theta^{\ast}$. Note that (15) is a discrete-time counterpart of the certificate in [25] for continuous-time systems. The proof of Theorem 2 is given in Appendix A. It is worth noting that the online safe control condition (17) requires only the estimated parameters without needing knowledge of $\theta^{\ast}$. This property makes (17) readily applicable for computing safe inputs via a safety filter, one that evaluates each nominal input and replaces it with the closest admissible input that renders the system safe (see Fig. 1(b)). Specifically, given a nominal state-feedback policy $\pi_{\mathrm{nom},t}:\mathcal{X}\to\mathbb{R}^{m}$ at time step $t$, a robust adaptive safe policy $\pi_{\mathrm{ras},t}:\mathcal{X}\to\mathcal{U}$ is implicitly given by:
$\pi_{\mathrm{ras},t}(x;p)=\operatorname*{arg\,min}_{\nu\in\mathcal{U}_{\mathrm{ras},t}(x,p)}\|\nu-\pi_{\mathrm{nom},t}(x)\|.\vskip-5.69046pt$ (20)
The safe policy $\pi_{\mathrm{ras},t}(\cdot;p)$ is time-varying partly because the underlying admissible set $\mathcal{U}_{\mathrm{ras},t}$ evolves over time, inherently capturing the influence of past states and inputs via the estimator $\mathtt{EST}$.
@@
The online input condition (17) is less conservative than the CBC (15), which is only used for offline design and verification of CBFs. Besides, the robustified barrier function (13) is built using the true parametric error $e_{\theta,t}$ instead of the set-induced error bound as in [25]. In fact, unlike what has been reported in [25], the invariance property of the safe sets defined by the maximized error bound-based robust adaptive CBFs can only be derived in discrete time, but not in continuous time. More details are given in Appendix B.
3.3 Practical implementation
Deriving a closed-form $\pi_{\mathrm{ras},t}(\cdot;p)$ is difficult in general. In practice, it suffices to compute $u_{\mathrm{ras},t}(p)=\pi_{\mathrm{ras},t}(x_{t};p)$ by solving (20) for the current state $x_{t}$. This relies on characterizing $\mathcal{U}_{\mathrm{ras},t}(x_{t};p)$ via the condition (15), which also depends on $\varepsilon_{\theta,t}(p)$. Computing $\varepsilon_{\theta,t}(2)$ generally requires solving a non-convex program, while using $\varepsilon_{\theta,t}(1)$ reduces (9b) to a linear program that can be solved efficiently. Since $\varepsilon_{\theta,t}(2)\leq\varepsilon_{\theta,t}(1)$, we have $\mathcal{U}_{\mathrm{ras},t}(x;1)\subseteq\mathcal{U}_{\mathrm{ras},t}(x;2)$, meaning the $1$-norm-based condition is more conservative but offers faster computation in practice.
In general, the optimization problem (20) is non-convex and cannot be solved efficiently. However, noting that $f$ is linear in $u$ due to the input-affine structure (cf. the model in (2.2)), when the designed barrier candidate $B$ is concave w.r.t. its first argument such that $\mathcal{U}_{\mathrm{ras},t}(x;p)$ is a convex set, then computing $u_{\mathrm{ras},t}(p)$ is equivalent to solving a convex optimization problem, for which many efficient algorithms with convergence and optimality guarantees exist [7]. Note that the concavity of $B$ w.r.t. the state $x$ is a standard assumption in many DT-CBF formulations [2], [44]; however, the design of concave model-based barrier functions depends on specific applications, which is out of the scope of the current paper.